OpenStudy (anonymous):

Another tricky problem.. @jim_thompson5910

OpenStudy (anonymous):

This is all I know: at 1st=1 at 2nd=1+5=6 at 3rd=6+7=13 at 4th=13+9=22 at 5th=22+11=33 at 6th=33+13=46 at 7th=46+15=61

jimthompson5910 (jim_thompson5910):

the 2nd term is when n = 2 notice we're adding 5 to the first term (1) to get the second term (6)

jimthompson5910 (jim_thompson5910):

so let's assume the answer is choice A that says "take the previous term and add n to it to get the next term" but n = 2 and we are not adding 2, we're adding 5

jimthompson5910 (jim_thompson5910):

so do you see how I eliminated choice A?

OpenStudy (anonymous):

I see, yes ! Progress is looking good! :)

jimthompson5910 (jim_thompson5910):

try out choice B

OpenStudy (anonymous):

How?

jimthompson5910 (jim_thompson5910):

notice we're now adding on 4n+1 to the previous term to get the next term

OpenStudy (anonymous):

N= 5 , correct?

jimthompson5910 (jim_thompson5910):

let's go back to choice A